OpenStudy (anonymous):

the principal is 800$ the rate is 8% for three months. what is the interest?

OpenStudy (jdoe0001):

\(\bf i = prt \\ \quad \\ p = 800\qquad r=8\%\to \frac{8}{100}\to 0.08\qquad t=3\ months\to \frac{1}{3}\ year\) plug in your values

OpenStudy (jdoe0001):

hmmm wait a sec.. shoot, 3months is not 1/3 of a year dohh anyhow \(\bf i = prt \\ \quad \\ p = 800\qquad r=8\%\to \frac{8}{100}\to 0.08\qquad t=3\ months\to \frac{3}{12}\ year\to \frac{1}{4}\)

OpenStudy (anonymous):

thanks for your help. I thought I was off.

OpenStudy (jdoe0001):

yw